# Do Generative AI Models Output Harm while Representing Non-Western Cultures: Evidence from A Community-Centered Approach

### v2

@@ 1 Introduction
Innovations into Generative Artificial Intelligence (GAI) tools such as text-to-image generators (T2Is) have revolutionized the way we create content, such as producing realistic images/videos. This potential has seen a sharp uptick in GAI adoption in industries such as healthcare (Hastings 2024) and policy-making (Huer and Saenz 2003), with the promise of unlocking unprecedented capabilities with far-reaching implications. However, GAI tools have also been found to embed biases that skew their generated artefacts towards the experiences of a select few rather than the diverse many (Septiandri et al. 2023; Gupta et al. 2023b), prompting further scrutiny of their potential effects on society.
One such pattern of bias is a ‘Western gaze,’ whereby outputs around non-Western contexts contain superimposed Western interpretations of those contexts (Kotliar 2020). This is a critical problem, especially given how T2Is are being used globally and beyond just the West, and perhaps most prominently in Global South contexts such as Asia and Africa (Bianchi et al. 2023; Venkit et al. 2023c; Gautam, Venkit, and Ghosh 2024). As representational systems (Hall 1997a), T2Is thus have the power to make meaning and influence culture through produced artefacts and, by applying a Western gaze to Non-Western contexts, cause ‘representational harms’ (Barocas et al. 2017) by misrepresenting and marginalizing the experiences of populations that are globally minoritized (Bender et al. 2021). However, this remains an under-addressed issue, as research into T2I ethics and fairness often center Western perspectives and rely on Western frameworks of ethics and fairness (Septiandri et al. 2023; Das et al. 2024; Venkit et al. 2023b).
Motivated by these gaps, our work is a community-centered investigation into depictions of non-Western cultures within T2I outputs, focusing on a specific context where AI-mediated harms are rising (Raj 2024): India, an incredibly diverse country with over 1.4 billion people spanning a wide range of regional subcultures, each with their own languages, cuisines, attire, art forms, festivals, and much more (Panda and Gupta 2004). Through grounded theory analysis (Charmaz 2017) of 5 focus groups with 25 participants from various Indian subcultures/regions, we explore the following questions:
@@
Beyond focus group participants perceiving pre-defined representational harms such as stereotyping of Indianness, the erasure and disparagement of specific subcultures and a lower quality of service in comparison to Western contexts (Dev et al. 2020), we document two novel cultural harms:
@@
    We also present the novel representational harm of cultural misappropriation within the outputs of T2Is, defined as “the depictions of details about a culturally-specific context which incorrectly embeds details from one culture or subculture outside of its appropriate context.” Some examples of cultural misappropriation observed within our focus groups include homogenizing commonly occurring Indian breakfast items from different parts of India onto the same plate, incorrectly presenting draping region-specific styles of wearing Indian clothing, and showing inaccurate ornaments for dance forms with centuries of historical and cultural contexts. In documenting this representational harm, we uncover how T2Is embed very surface-level understanding of details about Indian culture and its various subcultures.
Having documented such harms perceived by individuals from a wide range of Indian communities within live-generated outputs of the T2I Stable Diffusion in response to prompts of their own choosing, we present guidelines for design principles that can inform the development of more inclusive and culturally sensitive T2Is that can accurately represent nuanced cultural interactions and experiences. We adopt a sociotechnical lens to these principles (Shelby et al. 2022), recognizing the importance of culture-informed data work in alleviating these issues (Sambasivan et al. 2021b). We contribute to the large conversation around designing safer T2Is in non-Western contexts and the Global South (e.g., Gautam, Venkit, and Ghosh 2024; Qadri et al. 2023; Sambasivan et al. 2021b), as we consider the application of our proposed design principles in other non-Western cultures not centered in T2I design and suffering similar representational harms.
@@ 2 Related Work
2.1 Biases and Harms within GAI Tools: A Sociotechnical Systems Perspective
@@
The extensive documentation of biases embedded within GAI tools and the operationalization of such biases to produce harmful outcomes has sparked widespread concern among researchers, who have highlighted the far-reaching negative consequences of these models on diverse communities (Bender et al. 2021; Caliskan, Bryson, and Narayanan 2017; Gupta et al. 2023a). As such outputs are seen to amplify societal biases that disproportionately affect marginalized groups by perpetuating a variety of racial (Field et al. 2021; Sap et al. 2019), gender (Bolukbasi et al. 2016; Caliskan, Bryson, and Narayanan 2017), disability (Gadiraju et al. 2023; Venkit, Srinath, and Wilson 2023), and nationality-based biases (Venkit et al. 2023b, c) that reinforce outdated attitudes and exacerbating harmful stereotypes.
@@
Through the propagation of such stereotypes and the presence of negative biases, the outputs of GAI tools can and do cause harm. In the broader context of machine learning-based systems, (Blodgett 2021) break these down into representational and allocational harm, based on whether harmful representations are generalized or resources are distributed disparately. A taxonomy of fine-grained representational harm categories includes quality of service, stereotyping, stigmatization, alienation, and public participation. Documenting attributes of bias measures can facilitate a better understanding of harms, use cases, and limitations. Building on this, Dev et al. (2022) developed a framework categorizing harm into five types: Stereotyping, Disparagement, Dehumanization, Erasure, and Quality of Service. Their work offers a framework to GAI researchers such as ourselves for labeling and investigating patterns of harms.
@@
Investigating such harms has become especially critical as GAI systems are growing into a prominent part of the suite of AI-as-a-Service (AIaaS) tools that aim to provide convenient plug-and-play AI services (Lewicki et al. 2023) across interdisciplinary fields such as education (Rane et al. 2023), healthcare (Zack et al. 2023; Hastings 2024), and policy-making (Huer and Saenz 2003). The ease of using AIaaS tools has led to its meteoric rise in adoption in global contexts, and a subsequent surge in investigating how biases in such contexts propagate globally (e.g., Lewicki et al. 2023; Venkit 2023). While designers of GAI and AIaaS tools have been working on mitigating such biases, the majority of approaches have been found to be mostly technical (Birhane 2021), completely ignoring the social contexts in which systems operate (Venkit et al. 2023a, 2024).
@@
Researchers have therefore argued that mitigation of biases should adopt a sociotechnical approach: recognizing that AIaaS tools both influence and are impacted by social stakeholders and end-users (Cooper and Foster 1971; Brynjolfsson, Li, and Raymond 2023). This is important, since purely technical approaches miss the important fact that the data that drives GAI tools embedded within AIaaS services is heavily influenced by the perspectives and positionalities of the humans who were part of model training teams (Birhane 2021), and that such an oversight can lead to perceived technical ‘solutions’ being ineffective in the real world (D’ignazio and Klein 2020). A sociotechnical systems perspective on AI services and embedded GAI tools requires the perspectives of communities whose members experience the harms and suffer their impact in their daily lives. This approach prioritizes the equal involvement of social stakeholders and minority voices in the creation of these sociotechnical systems, ensuring that their perspectives and experiences are not erased or marginalized. Community-centered approaches towards recognizing and mitigating the biases within GAI tools, such as the work of Gadiraju et al. (2023), Qadri et al. (2023), Mack et al. (2024), and Prabhakaran, Qadri, and Hutchinson (2022), are thus essential.
@@
However, most research in understanding and mitigating harms has largely ignored such harms within non-Western contexts, with a few notable exceptions (Das et al. 2024; Qadri et al. 2023). Our study aims to address this gap by adopting a community-based approach and ethnographic principles to investigate and define harms generated by AI models, particularly emphasizing cultural and subculture representation within diverse communities across India.
2.2 Stereotypical Representations of Indian Culture/ Subcultures
@@
Though depictions of India in Western media as ‘the land of snake charmers’ (Chaudhuri 2009) and other features which Satyajit Ray (1958) called ‘the false-exotic’ might sound like relics of the past, that is not the case. As recently as the 2010 Julia Roberts classic ‘Eat, Pray, Love’, the protagonist perpetuates another famous Indian stereotype, as she visits India to ‘find herself.’ Hollywood projections of India are littered with stereotypes, both Orientalist perspectives (Said 1977) that show India as the “land of miracles and wonders,” (Chakkarath 2010) as well as ones about subcultures such as “the artistic Bengali, the simple Gujarati, the austere Maharashtrian, and the intelligent Tamil brahmin” (Mehta and Pandharipande 2011). Furthermore, given how films made in Hollywood or European film industries often are centered in their own contexts, Indian representation typically happens through characters in the Indian diaspora, either as fully Westernized individuals who ignore or actively look down upon elements of Indian culture (most commonly, eating with one’s hands), or projecting a romanticized view of India as a place they long to return to (Trivedi 2008). While portrayals of Indian culture within Western film media might evoke a sense of India having established a presence on the global stage (Matusitz and Payano 2012), stereotypical representations of Indian-ness still leave much to be desired.
@@
However, any discussion of representations of Indian culture/ subcultures is incomplete without addressing how they occur within Indian contexts. Mainstream media within India, primarily dominated by India’s Bollywood film industry, also perpetuates stereotypes around Indian culture/subcultures. As Arora and Sylvia (2023) writes, “popular Hindi cinema has articulated Indian identity as Hindu identity, with the normative Indian citizen presented as North Indian and Hindu.” Bollywood movies across the past 70 years has overwhelmingly featured characters with traditionally Hindu names, with little to no geographical representation of the Northeastern states of India (Khadilkar, KhudaBukhsh, and Mitchell 2022). Muslim characters, when shown, are often cast as violent antagonists or terrorists assosciated with Pakistan (Bhat 2019; Kumar 2013), or the ‘good Muslim’ (Mamdani 2002) who stands against other antagonistic Muslims (Islam 2007). Similarly, Northeast India, when featured, is often romanticized as exotic landscapes, contrasting the slow and rural lifestyles here with the busy, urban cities elsewhere in India (Dowerah and Nath 2017; Hasan 2011). These depictions of India contribute strongly to how Indian culture is perceived, defining ‘Indianness’ to the West (Matusitz and Payano 2011).
@@
Our paper directly follows Qadri et al. (2023)’s investigation into T2I representations of the Southeast Asian countries of India, Bangladesh and Pakistan. We extend this work by focusing specifically on Indian culture/subcultures, paying respect to their uniqueness within Southeast Asia and going against any suggestions of a regional monolith, and dive deeper into exploring how the depictions of cultures/subcultures actively cause harms to individuals who identify with such cultures/subcultures.
@@ 3.1 Recruitment and Focus Group Descriptions
We recruited participants through advertising the study through our social networks and postings in public places, inviting participation from individuals above the age of 18 and self-identifying as being Indian by virtue of growing up in India or not having lived outside of India for over 10 years. We successfully recruited 25 participants. Among these, 60% self-identified as female, while 40% self-identified as male. We asked participants which of the five regions of India (Northern, Eastern, Western, Southern, or Central) they self-identified as being from. Our study represents 4 participants from North India, 6 from East India, 5 from West India, 9 from South India, and 1 from Central India. We randomly sorted participants into five focus groups of 5 individuals each, and not grouping participants from similar regions to avoid groupthink bias where like-minded participants would only have shallow conversations (MacDougall and Baum 1997), with each focus group being conducted Oct-Dec 2023 by two researchers in accordance with best practices (Halcomb et al. 2007). We present participant distribution information in Table 1 (Appendix C), alongside assigned pseudonyms inspired by renowned literary figures from diverse cultural and subcultural backgrounds across India, to avoid numeric identifiers while still using culturally-appropriate names as opposed to Western pseudonyms commonly used. By using culturally meaningful pseudonyms, we avoid performing cultural erasure while respecting their anonymity as we highlight quotes throughout this paper.
@@
We asked participants to reflect on specific examples on areas they believed the images were to their satisfaction, as well as where they were dissatisfied, inviting them to discuss within the group rather than us asking each participant individually. We also offered participants the option of providing us with any ‘follow-up’ prompts for which they would like to see the outputs to, if such outputs could be used in comparison with those from the original prompts. Finally, we asked participants whether they believed that a given prompt output was an image they would show to non-Indian individuals as good representations of Indian culture or one or more subcultures, and why/why not. In support of open science and ensuring transparency in our research process, we have made the anonymized materials used for this study available in our repository [^1] https://github.com/PranavNV/Harms_In_GAI. This includes quotes from focus groups, coding frameworks, recruitment survey, generated images, and other data that can contribute to a better understanding of our work. The study received Exempt status by Institutional Review Board (IRB) review.
@@ 3.2 Thematic Analysis
We analyzed our focus group data through a constructivist grounded theory approach (Charmaz 2017), individually coding transcripts line-by-line in accordance with Charmaz (2006) to generate theories through inductive reasoning (Glaser and Strauss 1967; Glaser 1992). We conducted qualitative coding on the platform Taguette [^2] https://app.taguette.org/, creating codes and associated definitions for transcript snippets, and later synthesizing them through discussion within the research team. The codebook will be made available upon publication.

### v3

@@ 1 Introduction
Innovations into Generative Artificial Intelligence (GAI) tools such as text-to-image generators (T2Is) have revolutionized the way we create content, such as producing realistic images/videos. This progress has seen a sharp uptick in GAI adoption in industries such as healthcare (Hastings 2024) and policy-making (Huer and Saenz 2003), with the promise of unlocking unprecedented capabilities with far-reaching implications. However, GAI tools have also been found to embed biases that skew their generated artefacts towards the experiences of a select few rather than the diverse many (Septiandri et al. 2023; Gupta et al. 2023b), prompting further scrutiny of their potential effects on society.
One such pattern of bias is a ‘Western gaze,’ whereby outputs around non-Western contexts contain superimposed Western interpretations of those contexts (Kotliar 2020). This is a critical problem, especially given how T2Is are being used globally and beyond just the West, and perhaps most prominently in Global South contexts such as Asia and Africa (Bianchi et al. 2023; Gautam, Venkit, and Ghosh 2024). As representational systems (Hall 1997a), T2Is thus have the power to make meaning and influence culture through produced artefacts and, by applying a Western gaze to Non-Western contexts, cause ‘representational harms’ (Barocas et al. 2017) by misrepresenting and marginalizing the experiences of populations that are globally minoritized (Bender et al. 2021). However, this remains an under-addressed issue, as research into T2I ethics and fairness often center Western perspectives and rely on Western frameworks of ethics and fairness (Septiandri et al. 2023; Das et al. 2024; Venkit et al. 2023b).
We address these gaps through a Non-Western community centered investigation into depictions of non-Western cultures within T2I outputs, focusing on a specific context where AI-mediated harms are rising (Raj 2024): India, an incredibly diverse country with over 1.4 billion people across many regional subcultures, each with their own cuisines, attire, art forms, festivals, and much more (Panda and Gupta 2004). Through grounded theory analysis (Charmaz 2017) of 5 focus groups with 25 participants from various Indian subcultures/regions, we explore the following questions:
@@
Beyond focus group participants perceiving representational harms defined by Dev et al. (2020) such as stereotyping of Indianness, the erasure and disparagement of specific subcultures and a lower quality of service in comparison to Western contexts we document two novel cultural harms:
@@
    We also present the novel representational harm of cultural misappropriation within the outputs of T2Is, defined as “the depictions of details about a culturally-specific context which incorrectly embeds details from one culture or subculture outside of its appropriate context.” Some examples of cultural misappropriation observed within our focus groups include homogenizing commonly occurring Indian breakfast items from different parts of India onto the same plate, incorrectly presenting region-specific styles of wearing Indian clothing, and showing inaccurate ornaments for dance forms with centuries of historical and cultural contexts. In documenting this representational harm, we uncover how T2Is embed very surface-level understanding of details about Indian culture and its various subcultures.
Having documented such harms perceived by individuals from a wide range of Indian communities within live-generated outputs of the T2I Stable Diffusion in response to prompts of their own choosing, we present guidelines for design principles that can inform the development of more inclusive and culturally sensitive T2Is that can accurately represent nuanced cultural interactions and experiences. We adopt a sociotechnical lens (Shelby et al. 2022), recognizing the importance of culture-informed data work in alleviating these issues (Sambasivan et al. 2021b). We contribute to the large conversation around designing safer T2Is in non-Western contexts and the Global South (e.g., Gautam, Venkit, and Ghosh 2024; Qadri et al. 2023; Sambasivan et al. 2021b), as we consider the application of our proposed design principles in other non-Western cultures not centered in T2I design and suffering similar representational harms.
@@ 2 Related Work
2.1 Biases and Harms within GAI Tools:
       A Sociotechnical Systems Perspective
@@
The extensive documentation and operationalization of biases embedded within GAI tools to produce harmful outcomes has sparked widespread concern among researchers, who have highlighted the far-reaching negative consequences of these models on diverse communities (Bianchi et al. 2023; Bender et al. 2021; Gupta et al. 2023a). As such outputs are seen to amplify societal biases that disproportionately affect marginalized groups by perpetuating a variety of racial (Field et al. 2021), gender (Bolukbasi et al. 2016; Caliskan, Bryson, and Narayanan 2017), disability (Gadiraju et al. 2023; Venkit, Srinath, and Wilson 2023), and nationality-based biases (Venkit et al. 2023b, c) that reinforce outdated attitudes and exacerbating harmful stereotypes.
@@
Through the propagation of such stereotypes and the presence of negative biases, the outputs of GAI tools can and do cause harm (Das et al. 2024). In the broader context of machine learning-based systems, Blodgett (2021) break these down into representational and allocational harm, based on whether harmful representations are generalized or resources are distributed disparately. A taxonomy of fine-grained representational harm categories includes quality of service, stereotyping, stigmatization, alienation, and public participation (Blodgett 2021). Documenting attributes of bias measures can facilitate a better understanding of harms, use cases, and limitations. Building on this, Dev et al. (2022) developed a framework categorizing harm into five types: Stereotyping, Disparagement, Dehumanization, Erasure, and Quality of Service. Their work offers a framework to GAI researchers such as ourselves for labeling and investigating patterns of harms.
@@
Investigating such harms has become especially critical as GAI systems are growing into a prominent part of the suite of AI-as-a-Service (AIaaS) tools that aim to provide convenient plug-and-play AI services (Lewicki et al. 2023) across broad domains such as education (Rane et al. 2023), healthcare (Zack et al. 2023; Hastings 2024), and policy-making (Huer and Saenz 2003). The ease of using AIaaS tools has led to its meteoric rise in adoption in global contexts, and a subsequent surge in investigating how biases in such contexts propagate globally (e.g., Lewicki et al. 2023; Venkit 2023). While designers of GAI and AIaaS tools have been working on mitigating such biases, the majority of approaches have been found to be mostly technical (Birhane 2021), completely ignoring the social contexts in which systems operate (Venkit et al. 2023a, 2024).
@@
Researchers have therefore argued that mitigation of biases should adopt a sociotechnical approach, recognizing that AIaaS tools both influence and are impacted by end-users (Cooper and Foster 1971; Brynjolfsson, Li, and Raymond 2023). This is critical since purely technical approaches miss the important fact that the data driving GAI tools embedded within AIaaS services is heavily influenced by the perspectives and positionalities of the humans who were part of model training teams (Birhane 2021), and that such an oversight can lead to perceived technical ‘solutions’ being ineffective in the real world (D’ignazio and Klein 2020). A sociotechnical systems perspective on AI services and embedded GAI tools requires the perspectives of communities whose members experience the harms and suffer their impact in their daily lives. This approach prioritizes the equal involvement of social stakeholders and minority voices in the design, development and deployment of these sociotechnical systems, ensuring that their perspectives and experiences are not erased or marginalized. Community-centered approaches towards recognizing and mitigating the biases within GAI tools, such as the work of Qadri et al. (2023), Mack et al. (2024), and Prabhakaran, Qadri, and Hutchinson (2022), are thus essential.
@@
However, most research in understanding and mitigating harms has largely ignored such harms within non-Western contexts, with a few notable exceptions (Das et al. 2024; Qadri et al. 2023). Our study aims to address this gap by adopting a community-based approach to investigate and define harms generated by AI models, particularly emphasizing cultural and subcultural representation within diverse communities across India.
2.2 Stereotypical Representations of
       Indian Culture/ Subcultures
@@ 2.2 Stereotypical Representations of
Though depictions of India in Western media as ‘the land of snake charmers’ (Chaudhuri 2009) and other features which Satyajit Ray (1958) called ‘the false-exotic’ might sound like relics of the past, that is not the case. As recently as the 2010 Julia Roberts classic ‘Eat, Pray, Love,’ the protagonist perpetuates another famous Indian stereotype, as she visits India to ‘find herself.’ Hollywood projections of India are littered with stereotypes, both Orientalist perspectives (Said 1977) that show India as the “land of miracles and wonders,” (Chakkarath 2010) as well as ones about subcultures such as “the artistic Bengali, the simple Gujarati, the austere Maharashtrian, and the intelligent Tamil brahmin” (Mehta and Pandharipande 2011). Furthermore, given how films made in Hollywood or European film industries often are centered in their own contexts, Indian representation typically happens through characters in the Indian diaspora, either as fully Westernized individuals who ignore or actively look down upon elements of Indian culture (most commonly, eating with one’s hands), or projecting a romanticized view of India as a place they long to return to (Trivedi 2008). While portrayals of Indian culture within Western film media might evoke a sense of India having established a presence on the global stage (Matusitz and Payano 2012), stereotypical representations of Indianness still leave much to be desired.
@@
However, any discussion of representations of Indian culture/subcultures is incomplete without addressing how they occur within Indian contexts. Mainstream media within India, primarily dominated by India’s Bollywood film industry, also perpetuates stereotypes around Indian culture/subcultures. As Arora and Sylvia (2023) writes, “popular Hindi cinema has articulated Indian identity as Hindu identity, with the normative Indian citizen presented as North Indian and Hindu.” Bollywood movies across the past 70 years has overwhelmingly featured characters with traditionally Hindu names, with little to no geographical representation of the Northeastern states of India (Khadilkar, KhudaBukhsh, and Mitchell 2022). Muslim characters, when shown, are often cast as violent antagonists or terrorists associated with Pakistan (Bhat 2019; Kumar 2013), or the ‘good Muslim’ (Mamdani 2002) who stands against other antagonistic Muslims (Islam 2007). Similarly, Northeast India, when featured, is often romanticized as exotic landscapes, contrasting the slow and rural lifestyles here with the busy, urban cities elsewhere in India (Dowerah and Nath 2017; Hasan 2011). These depictions of India contribute strongly to how Indian culture is perceived, defining ‘Indianness’ to the West (Matusitz and Payano 2011).
@@
Our paper directly follows Qadri et al. (2023)’s investigation into T2I representations of the Southeast Asian countries of India, Bangladesh and Pakistan. We extend this work by focusing specifically on Indian culture/subcultures, paying respect to their uniqueness within Southeast Asia and going against any suggestions of a regional monolith, and dive deeper into showing how the depictions of cultures/subcultures actively cause harms to individuals who identify with such cultures/subcultures.
@@ 3.1 Recruitment and Focus Group Descriptions
We recruited participants through advertising the study through our social networks and postings in public places, inviting participation from individuals above the age of 18 and self-identifying as being Indian by virtue of growing up in India or not having lived outside of India for over 10 years. We successfully recruited 25 participants. Among these, 60% self-identified as female, while 40% self-identified as male. We asked participants which of the five regions of India (Northern, Eastern, Western, Southern, or Central) they self-identified as being from. Our study represents 4 participants from North India, 6 from East India, 5 from West India, 9 from South India, and 1 from Central India. We randomly sorted participants into 5 focus groups of 5 individuals each, and not grouping participants from similar regions to avoid groupthink bias where like-minded participants would only have shallow conversations (MacDougall and Baum 1997), with each focus group being conducted Oct-Dec 2023 by two researchers in accordance with best practices (Halcomb et al. 2007). We present participant distribution information in Table 1 (Appendix C), alongside assigned pseudonyms inspired by renowned literary figures from diverse cultural and subcultural backgrounds across India, to avoid numeric identifiers while still using culturally-appropriate names as opposed to Western pseudonyms commonly used. By using culturally meaningful pseudonyms, we avoid performing cultural erasure while respecting their anonymity as we highlight quotes throughout this paper.
@@
We asked participants to reflect on specific examples on areas they believed the images were to their satisfaction, as well as where they were dissatisfied, inviting them to discuss within the group rather than us asking each participant individually. We also offered participants the option of providing us with any ‘follow-up’ prompts for which they would like to see the outputs to, if such outputs could be used in comparison with those from the original prompts. Finally, we asked participants whether they believed that a given prompt output was an image they would show to non-Indian individuals as good representations of Indian culture or one or more subcultures, and why/why not. In support of open science and ensuring transparency in our research process, we have made the anonymized materials used for this study available in our repository [^1] https://github.com/PranavNV/Harms˙In˙GAI. This includes quotes from focus groups, coding frameworks, recruitment survey, generated images, and other data that can contribute to a better understanding of our work. The study received Exempt status by Institutional Review Board (IRB) review.
@@ 3.2 Thematic Analysis
We analyzed our focus group data through a constructivist grounded theory approach (Charmaz 2017), individually coding transcripts line-by-line in accordance with Charmaz (2006) to generate theories through inductive reasoning (Glaser and Strauss 1967; Glaser 1992). We conducted qualitative coding on the platform Taguette [^2] https://app.taguette.org/, creating codes and associated definitions for transcript snippets, and later synthesizing them through discussion within the research team.